Secret:

Give the values of a, b, and c needed to write the equation's standard form. (5 + x)(5 - x) = 7 A = 1; B = 0; C = -18 A = 25; B = 0; C = -1 A = -1; B = 0; C = 25

jhonyy9:

do you know what is the standard form of an equation ?

jhonyy9:

ax +by =c

jhonyy9:

do you know formula of difference of squares ?

MARC:

Standard eqn form means---> \(ax^2+bx=-c\)

Secret:

Ohhh

MARC:

Step 1 : Expand this first \((5+x)(5-x)=7\)

MARC:

u will get \(25+5x-5x-x^2=7\)

MARC:

Step 2 : Rearrange in standard eqn form \(ax^2+bx=-c\)

MARC:

\(x^2+0x=18\)

MARC:

Step 3 : Compare the values to find a,b and c

MARC:

\(\color{red}{1}x^2+\color{blue}{0}x=\color{green}{18}\) compare with \(\color{red}{a}x^2+\color{blue}{b}x=\color{green}{-c}\)

MARC:

Therefore, \(a=1\) \(b=0\) \(c=-18\)

Secret:

Ohhh

MARC:

There is another method to find the correct values of A,B and C

MARC:

Just simply substitute the value of A,B and C into this eqn \(Ax^2+Bx=-C\)

MARC:

If the eqn is equal to \(1x^2+0x=-18\) Therefore,that is the answer

Secret:

Oof, ok Thanks!

MARC:

Yw! ^